# A continuous symmetry breaking measure for finite clusters using Jensen-Shannon divergence

### v1

@@ 7 Conclusion
In this paper, we introduce a continuous SBM using the Jensen-Shannon divergence to analyze structural transformations and distortions. It is designed for studying SB in finite clusters where SB is continuous and there is value in quantifying it. In crystals, a symmetry is either present or broken, and it cannot be slightly broken. The continuous SBM, on the contrary, provides insights into local SB that would otherwise be obscured in the analysis of the infinite crystal.
@@ Acknowledgment
Supplemental Materials
Section 3.2.2
Let $F$ and $G$ be two three-dimensional normal variables with means separated by vector $\bm{d}$ and identical covariance matrices $U\bm{I}_{3}$, where $U>0$ is a constant. Their probability density functions are:
$f(\bm{x})=\mathcal{N}(\bm{x};\bm{\mu},U\bm{I}_{3}),\quad g(\bm{x})=\mathcal{N}(\bm{x};\bm{\mu}+\bm{d},U\bm{I}_{3}).$ (18)
KL Divergence Calculation
The Kullback-Leibler divergence for two $n$-dimensional normal distributions $u$ and $v$ with means $\bm{\mu}^{u}$ and $\bm{\mu}^{v}$, and covariance matrices $\bm{\Sigma}^{u}$ and $\bm{\Sigma}^{v}$ respectively, is given by [19]:
$D_{\mathrm{KL}}(u\|v)=\frac{1}{2}\log\frac{\left|\bm{\Sigma}^{v}\right|}{\left|\bm{\Sigma}^{u}\right|}+\frac{1}{2}\operatorname{Tr}\left((\bm{\Sigma}^{v})^{-1}\bm{\Sigma}^{u}\right)+\frac{1}{2}\left(\bm{\mu}^{u}-\bm{\mu}^{v}\right)^{\top}\left(\bm{\Sigma}^{v}\right)^{-1}\left(\bm{\mu}^{u}-\bm{\mu}^{v}\right)-\frac{n}{2}.$ (19)
For our case, since the covariance matrices are identical and equal to $U\bm{I}_{3}$, the formula simplifies as follows:
$D_{\mathrm{KL}}(f\|g)=\frac{1}{2}\operatorname{Tr}(\bm{I}_{3})+\frac{1}{2U}|\bm{d}|^{2}-\frac{3}{2}=\frac{d^{2}}{2U},$ (20)
@@
Section 5.1.2
Deviation from four-fold rotational symmetry $C_{4}$
SBM with displacements on corner atoms
Figure 8: The SBM $\mathscr{S}^{JS}_{C_{4}}[\mu_{d}]$ of a cubic cutout from the distorted Nickel structure. One of its corner atoms (darker blue atoms) is shifted parallel to the a-b plane (blue), deviated from the a-b plane by 45 degrees (red), or perpendicular to the a-b plane (yellow). The SBM is calculated for the four-fold rotational operation $C_{4}$, whose axis passes through the center of the finite cluster, with its direction along the positive c-axis. The figure is plotted as a function of the length of the atom displacement $d$.
In Figure 8, the displacement dynamics of one of the corner atoms within the nickel unit cell are examined. When the corner atom moves in a direction that does not interfere with other atoms in the lattice, the SBM $\mathscr{S}^{JS}_{C_{4}}[\mu_{d}]$ rapidly converges to $0.143$ ($\approx 2/14$). This convergence value matches that observed for face atoms, suggesting a uniform threshold in the SBM associated with deviations from a $C_{4}$ rotational symmetry in the context of a single atom displacement within this nickel unit cell. Additionally, a local minimum at $d=2.49$ is identified, which is indicative of a scenario where a corner atom overlaps with its adjacent face atom. Further analysis anticipates the presence of additional local minima at $d=3.52$, $d=\sqrt{2}\times 3.52\approx 4.98$, and $d=\sqrt{3}\times 3.52\approx 6.10$, which correspond to other significant lattice configurations, where the displaced corner atom aligns with other critical points in the unit cell.
Deviation from the reflection operator $\sigma_{h}$
SBM with displacements on out of plane atoms
Figure 9: The SBM $\mathscr{S}^{JS}_{\sigma_{h}}[\mu_{d}]$ of a cubic cutout from the Nickel structure when one of its out-of-plane atoms (darker blue atoms) is shifted parallel to the plane (blue), deviated from the plane by 45 degrees (red), or perpendicular to the plane (yellow). The SBM is calculated for the reflection operation $\sigma_{h}$, whose mirror plane passes through the center of the finite cluster, with its normal vector along the positive $z$-axis. The figure is plotted as a function of the length of the atom displacement $d$.
We consider the more general case of out-of-plane atoms. As illustrated in Fig. 9, the response of $\mathscr{S}^{JS}_{\sigma_{h}}[\mu]$ to atomic displacements oriented parallel, at a 45-degree angle, and perpendicular to the mirror plane coincide for small $d$. The $\mathscr{S}^{JS}_{\sigma_{h}}[\mu]$ corresponding to parallel distortion (depicted in yellow) achieves a local minimum of $0.072$ at $d=1.76$. This coincides with scenarios where one of the out-of-plane atoms intersects the reflection plane. For the $\mathscr{S}^{JS}_{\sigma_{h}}[\mu]$ subjected to a 45-degree angle distortion (represented in red), a local minimum of $0.072$ is observed at $d=2.49$. This value is approximately $\frac{\sqrt{2}}{2}\times 3.52$$\approx 2.49$. This scenario mirrors the former, where a corner atom intersects the reflection plane. Lastly, the $\mathscr{S}^{JS}_{\sigma_{h}}[\mu]$ with a parallel angle distortion (depicted in blue) attains its local minimum at a proximate $d=2.49$, registering a value of $0.091$. This is indicative of a situation wherein one of the center out-of-plane atoms overlaps with one of its adjacent corner atoms.
Sample size analysis
We make a note here to highlight certain numerical intricacies encountered during our tests. As discussed, we employ a Monte Carlo simulation to estimate the SBM. We found that the sample size required to obtain an estimate of the SBM at a certain level of accuracy increases as $T_{\alpha}[\mu]$ deviates further from $\mu$. This implies that as atomic displacements magnify, Monte Carlo calculations demand more extensive sample sizes to yield estimates of SBM with consistent precision. In Fig. 10, we plot the sample sizes we used to estimate $\mathscr{S}^{JS}_{\sigma_{h}}[\mu]$. These sizes were determined considering a $95\%$ confidence interval and a bilateral error tolerance of $0.0025$. The procedure for computing the sample size is delineated in Algorithm 2. All the simulation tests addressed in this subsection employ a sample size determined through this methodology.
Figure 10: The recommended sample size for Monte Carlo simulations of $\mathscr{S}^{JS}_{\sigma_{h}}[\mu]$. The sample size for (a) finite clusters with displacements of in-plane atoms and (b) for displacements of out-of-plane atoms. The colors of the curves indicate the sample size for displacements parallel to the plane (blue), 45 degrees to the plane (red), and perpendicular to the plane (yellow). These were the sample sizes needed to give $\mathscr{S}^{JS}_{\sigma_{h}}[\mu]$ with a $95\%$ confidence interval and an error tolerance of .0025.0025.0025.0025 on each side.

### v2

@@ VII Conclusion
In this paper, we introduce a continuous SBM using the Jensen-Shannon divergence to analyze structural transformations and distortions. It is designed for studying SB in finite clusters where SB is continuous and there is value in quantifying it. In infinite crystals, a symmetry is either present or broken, and it cannot be slightly broken. The continuous SBM, on the contrary, provides insights into local SB that would otherwise be obscured in the analysis of the infinite crystal.
@@ Acknowledgment
Appendix A Appendix A: Derivation of Symmetry Breaking Measure for a Single-atom System
Let $\mu(\mathbf{x})=\mathcal{N}(\mathbf{0},U\cdot\bm{I}_{3})$ and $(T_{\bm{d}})_{\#}\mu(\mathbf{x})=\mathcal{N}(\mathbf{d},U\cdot\bm{I}_{3})$. Writing the densities,
$\mu(\mathbf{x})=(2\pi U)^{-3/2}\exp\!\Big(-\tfrac{1}{2U}\|\mathbf{x}\|^{2}\Big),\quad(T_{\bm{d}})_{\#}\mu(\mathbf{x})=(2\pi U)^{-3/2}\exp\!\Big(-\tfrac{1}{2U}\|\mathbf{x}-\mathbf{d}\|^{2}\Big).$ (21)
Since the normalizing constants coincide, the log-ratio simplifies to
$\log\frac{\mu(\mathbf{x})}{(T_{\bm{d}})_{\#}\mu(\mathbf{x})}=-\frac{1}{2U}\|\mathbf{x}\|^{2}+\frac{1}{2U}\|\mathbf{x}-\mathbf{d}\|^{2}=\frac{1}{2U}\!\left(\|\mathbf{x}-\mathbf{d}\|^{2}-\|\mathbf{x}\|^{2}\right)=\frac{1}{2U}\left(\|\mathbf{d}\|^{2}-2\,\mathbf{x}^{\top}\mathbf{d}\right).$ (22)
Therefore,
$D_{\mathrm{KL}}(\mu\|(T_{\bm{d}})_{\#}\mu)=\mathbb{E}_{\mu}\!\left[\log\frac{\mu(\mathbf{x})}{(T_{\bm{d}})_{\#}\mu(\mathbf{x})}\right]=\frac{1}{2U}\left(\|\mathbf{d}\|^{2}-2\,\mathbf{d}^{\top}\mathbb{E}_{\mu}[\mathbf{x}]\right).$
Because $\mu$ is centered and spherically symmetric, $\mathbb{E}_{\mu}[\mathbf{x}]=\mathbf{0}$. Thus, the linear cross term vanishes and we obtain the expression
$D_{\mathrm{KL}}(\mu\|(T_{\bm{d}})_{\#}\mu)=\frac{\|\mathbf{d}\|^{2}}{2U}=\frac{d^{2}}{2U},$